A hot air balloon can contain upwards of 16,000 L of gas (at STP)

how many combined molecules/atoms of air will be present?


1
Expert's answer
2022-03-20T09:55:12-0400

"n=\frac{V}{Vm}=\frac{16000}{22.4}=714.29" mol


"N=N(A)\times n=714.29\times6\times10^{23}=4285.71\times10^{23}"